# Lightweight Session-Key Rekeying Framework for Secure IoT–Edge Communication

### v1

@@ Abstract
The proliferation of Internet of Things (IoT) networks demands security mechanisms that protect constrained devices without the computational cost of public-key cryptography. Conventional Pre-Shared Key (PSK) encryption, while efficient, remains vulnerable due to static key reuse, replay attacks, and the lack of forward secrecy. This paper presents the Dynamic Session Enhanced Key Protocol (DSEKP)—a lightweight session-key rekeying framework, a fully symmetric extension to PSK that derives per-session AES–GCM keys using the HMAC-based Key Derivation Function (HKDF–SHA256) and authenticates session establishment through an HMAC proof in a single init–ack exchange. DSEKP was implemented on an ESP32 IoT sensor node and a Raspberry Pi 5 edge server communicating through a Mosquitto MQTT broker, and benchmarked against a static PSK baseline over more than 6,500 encrypted packets per configuration. The results demonstrate nearly identical throughput and reliability, with moderate overhead—mean latency increased by 27% and payload size by 10%—while delivering per-session forward secrecy and built-in replay protection. These findings confirm that dynamic symmetric rekeying can substantially strengthen IoT–Edge links with minimal computational and bandwidth cost, offering a practical migration path from static PSK to session-aware, scalable, and reproducible IoT security.
@@ 1 Introduction
IoT nodes such as microcontrollers and sensors operate with limited memory, processing power, and energy resources, rendering traditional public-key cryptographic protocols—such as Transport Layer Security (TLS) or Datagram TLS (DTLS) non-efficient [1] for large-scale deployment. Consequently, symmetric encryption based on the Pre–Shared Key (PSK) model remains a prevalent choice for securing lightweight protocols like MQTT and CoAP  [2, 3] due to its simplicity, low computational cost, and deterministic timing behavior.
Despite its efficiency, static PSK encryption exhibits fundamental weaknesses. The reuse of a single long-term key across all sessions compromises both past and future communications once the key is exposed [4]. Moreover, static PSK lacks forward secrecy, enabling replay and impersonation attacks. The operational challenges of securely distributing, rotating, and updating PSKs further hinder scalability in dynamic IoT networks [5].
To overcome these limitations, this study introduces the Dynamic Session Enhanced Key Protocol (DSEKP)—a lightweight, symmetric-only enhancement to the PSK framework that employs the HMAC-based Key Derivation Function (HKDF–SHA256) to generate ephemeral per-session encryption keys. DSEKP performs a single-round initialization handshake authenticated by an HMAC proof (InitProof) and derives fresh AES–GCM session keys from a combination of a device nonce, session counter, and timestamp. This design retains the simplicity of PSK while achieving forward secrecy, replay protection, and dynamic key refresh without relying on public-key cryptography or certificate infrastructures.
@@
    Lightweight symmetric session protocol: DSEKP extends PSK with HKDF-based per-session key derivation and HMAC authentication, achieving forward secrecy and replay protection without public-key operations.
@@ 2 Related Work
The security of Internet of Things (IoT) communication remains an active research field focused on balancing computational efficiency with cryptographic robustness. While asymmetric standards such as TLS 1.3 and DTLS 1.3 provide confidentiality and mutual authentication through public-key exchanges, their multi-round handshakes and certificate management overheads management are challenging for constrained microcontrollers such as the ESP32, ATmega, or STM32 families [6, 7]. Consequently, lightweight symmetric models based on the Pre–Shared Key (PSK) paradigm have been widely adopted for protocols including MQTT and CoAP [8, 9]. This section traces the evolution from static PSK schemes toward dynamic or session-based key management approaches, emphasizing the gaps that motivate the proposed DSEKP framework.
@@ 2.1 PSK-Based Encryption in IoT Systems
PSK architectures rely on a single long-term symmetric key shared between device and broker to perform AES–GCM encryption with minimal processing overhead. Studies such as  [10],  [11] and  [12] have shown efficient resource utilization on ESP32 microcontrollers using PSK–AES modes. However, static key reuse introduces significant vulnerabilities: once the PSK is compromised, all past and future traffic becomes decryptable, violating forward secrecy. Replay protection in such systems typically depends on timestamp or counter validation, which can desynchronize during Wi–Fi disruptions or low-power sleep cycles [13, 14]. Moreover, at scale, managing thousands of PSKs manually increases operational burden and risks inconsistent key distribution, leaving gaps exploitable by adversaries [15].
@@ 2.2 Dynamic Session-Key and Rekeying Protocols
To mitigate static-key weaknesses, researchers have explored session-based and hybrid key-management protocols. Protocols such as DTLS 1.3 and EDHOC (Ephemeral Diffie–Hellman Over COSE) [16, 17] achieve forward secrecy through elliptic-curve key exchanges, but these computations are prohibitively expensive for constrained MCUs [18, 19, 20]. In contrast, symmetric-only methods refresh session keys using derivation functions without public-key operations. The HMAC-based Key Derivation Function (HKDF) is particularly suited for IoT systems, as it expands a static secret using per-session entropy sources [21, 22, 23]. For instance,  [24, 25] employed HKDF–SHA256 for MQTT rekeying but required a centralized key server to manage nonces.  [26, 27, 28] shown hash-chain–based key evolution to enhance secrecy, yet synchronization overhead limited scalability. Efforts such as DTLS [29, 30] reduce handshake size through message compression but still depend on asymmetric initialization, leaving open the need for a purely symmetric, session-oriented protocol.
@@ 2.3 Comparative Gaps and Motivation
Comparative analyses consistently reveal a trade-off between cryptographic strength and computational feasibility.  [31] reported that ECDHE-based TLS increased energy consumption relative to PSK. Purely symmetric PSK models remain efficient but lack key agility and entropy mixing, making them susceptible to replay and traffic-analysis attacks [32, 33]. Furthermore, existing HKDF-based schemes often depend on centralized infrastructures or omit practical hardware validation. Hence, a clear gap exists for a fully symmetric, self-contained, and experimentally validated session-key protocol that offers forward secrecy, replay protection, and low overhead—without reliance on public-key infrastructures (PKI).
@@ 2.4 Contribution Context
The proposed Dynamic Session Enhanced Key Protocol (DSEKP) directly addresses these limitations. It preserves PSK’s lightweight efficiency while introducing HKDF-derived per-session keys and an HMAC-authenticated initialization handshake. Unlike prior HKDF or hash-chain approaches, DSEKP performs all derivation and verification steps locally on constrained devices without reliance on external key servers. Thus, it bridges the gap between the simplicity of static PSK and the forward secrecy of TLS, delivering a balanced, reproducible, and deployable solution for secure IoT–Edge communication.
@@ 3 Methodology and Framework
The proposed Dynamic Session Enhanced Key Protocol (DSEKP) extends the conventional Pre–Shared Key (PSK) model by introducing per-session dynamic key derivation and symmetric-only authentication using the HMAC-based Key Derivation Function (HKDF–SHA256). The protocol preserves the lightweight characteristics of PSK but replaces the static, reused key with ephemeral AES–GCM session keys derived from fresh entropy sources—device nonce, session counter, and timestamp—thereby ensuring forward secrecy and replay protection. This section describes the system architecture, data communication workflow, protocol phases, algorithmic details, and integrated security mechanisms.
@@ 3.3 Protocol Workflow
    Session Termination / Rotation: Upon device reboot or timeout, a new session counter triggers a fresh HKDF derivation, ensuring forward secrecy and automatic session cleanup on the edge node.
@@ 3.5 Security Operations
Forward Secrecy HKDF Entropy Mixing Ensures each session key is independent of previous sessions.
@@ 3.7 Summary
The DSEKP framework transforms the conventional static PSK architecture into a session-adaptive, stateless, and authenticated communication model for constrained IoT devices. By combining HKDF-based entropy mixing with HMAC-driven session authentication, DSEKP achieves TLS-grade confidentiality and forward secrecy with negligible additional cost in latency or bandwidth.
@@ 5.1 Descriptive Statistics
Interpretation and Justification: The results in Table 6 are derived directly from synchronized send–receive timestamps collected across 6,500 packet transmissions for each protocol, ensuring high statistical confidence. The 95 % confidence intervals (CI) confirm narrow uncertainty margins due to the large sample size and consistent network conditions. The observed 27 % increase in mean latency under DSEKP originates from the additional HKDF–SHA256 key derivation and HMAC verification performed once per session, introducing a small computational delay without affecting throughput stability. Similarly, the 10 % growth in mean payload corresponds to the inclusion of session metadata—namely the 2–byte session counter and 14–byte HMAC proof—required for forward secrecy and authentication. Despite these expected overheads, both throughput and reliability remain nearly identical between PSK and DSEKP, confirming that the proposed dynamic rekeying mechanism introduces negligible runtime or bandwidth penalty. These results verify that DSEKP preserves the lightweight characteristics of PSK while enhancing security and timing determinism in constrained IoT environments.
@@ 5.4 Payload and Overhead Analysis
Figure 7: Payload size comparison between PSK and DSEKP. The $\approx$ 10 % increase corresponds to additional session–metadata fields required for dynamic authentication. This minor overhead enables forward secrecy and replay protection without significant bandwidth penalty.
@@ 5.5 Security–Performance Trade–off Summary
Interpretation. DSEKP exhibits a moderate latency increase of approximately 27 % and a payload growth of about 10 % compared with PSK, while preserving identical throughput (1 pps) and near–perfect reliability (>>99.8 %). Although the latency difference is statistically significant (very low pp–values), the effect size remains moderate (Cohen’s $d\approx-0.49$), confirming that the overhead is small and within operational limits for lightweight IoT telemetry. Overall, DSEKP strengthens cryptographic resilience—achieving forward secrecy and replay protection—with negligible runtime penalty, offering a practical balance between performance and security for IoT–Edge deployments.
@@ 5.7 Interpretation of Results
The overall findings confirm that DSEKP effectively strengthens PSK–based security while preserving the lightweight characteristics required in IoT networks. The modest latency increase remains well within acceptable bounds for low–frequency telemetry ($\leq$ 1 Hz). DSEKP’s symmetric and stateless architecture allows seamless deployment on resource–constrained microcontrollers and edge gateways without modifying existing MQTT infrastructures. In summary, DSEKP delivers modern cryptographic agility—forward secrecy, replay protection, and session isolation—at negligible computational and bandwidth cost, establishing it as a practical upgrade path for PSK systems in real IoT–Edge environments.
@@ 6.1 Security Improvements Over PSK
DSEKP derives a unique AES–GCM key for every session using the HMAC-based Key Derivation Function (HKDF–SHA256), mixing device entropy sources—a 12-byte random nonce (DevNonce), 2-byte session counter (SessCtr), and 4-byte timestamp (TT)—with a long-term secret and salt. This ensures that compromise of one session key cannot expose previous or future sessions, thereby achieving practical forward secrecy without asymmetric cryptography.
@@ 6.2 Performance and Scalability Considerations
Despite introducing dynamic key derivation and per-session authentication, DSEKP’s runtime performance remains near-identical to static PSK. As shown in Figures 46, the mean latency increase of 27 % and 10 % payload growth are modest relative to the gains in forward secrecy and replay protection, remaining within typical Wi–Fi variance for low-frequency telemetry (1–2 Hz).
@@ 6.3 Comparative Perspective
When positioned against standard IoT security frameworks such as DTLS 1.3, EDHOC, or LAKE, DSEKP provides a pragmatic middle ground between performance and cryptographic assurance. It delivers TLS-grade forward secrecy using only symmetric primitives—achieving more than 80 % lower computational cost than DTLS while maintaining comparable latency to PSK.
@@ 6.6 Summary of Discussion
Overall, DSEKP transforms static PSK schemes into secure, adaptive, and stateless architectures suitable for modern IoT–Edge ecosystems. It provides measurable improvements in confidentiality, integrity, and forward secrecy while maintaining sub-second latency and constant throughput. By combining HKDF-based entropy mixing with symmetric HMAC authentication, DSEKP represents a practical and scalable step toward trustworthy, autonomous, and energy-efficient IoT communication.
@@ 7 Conclusion and Future Work
This paper presented the Dynamic Session Enhanced Key Protocol (DSEKP), a lightweight and symmetric-only security framework that strengthens conventional Pre–Shared Key (PSK) encryption for Internet of Things (IoT) edge environments. By deriving fresh AES–GCM session keys through HKDF–SHA256 and authenticating session initialization via an HMAC-based proof, DSEKP achieves per-session forward secrecy, replay protection, and stateless key management—without relying on public-key cryptography.
@@
The proposed protocol effectively bridges the gap between PSK simplicity and TLS-grade confidentiality, providing a deployable and scalable security solution for IoT–Edge communication. Its stateless, symmetric design supports seamless scaling to thousands of devices while maintaining strong per-session cryptographic isolation. In doing so, DSEKP establishes a practical foundation for trust-based and distributed edge security architectures in next-generation networks.
@@ Acknowledgments
Preprint submitted to arXiv (November 2025). The data and source codes will be made available upon reasonable request.

### v2

@@ Abstract
[…]
@@ 1 Introduction
IoT nodes such as microcontrollers and sensors operate with limited memory, processing power, and energy resources, rendering traditional public-key cryptographic protocols—such as Transport Layer Security (TLS) or Datagram TLS (DTLS) non-efficient [han2015semantic] for large-scale deployment. Consequently, symmetric encryption based on the Pre–Shared Key (PSK) model remains a prevalent choice for securing lightweight protocols like MQTT and CoAP  [shyam2025survey, bideh2020energy] due to its simplicity, low computational cost, and deterministic timing behavior.
Despite its efficiency, static PSK encryption exhibits fundamental weaknesses. The reuse of a single long-term key across all sessions compromises both past and future communications once the key is exposed [kuo2006comparison]. Moreover, static PSK reuses a single long-term key, exposing all past and future encrypted traffic if the key is compromised, and lacks replay protection when no freshness mechanisms are applied. The operational challenges of securely distributing, rotating, and updating PSKs further hinder scalability in dynamic IoT networks [michaels2018high].
To overcome these limitations, this study introduces the Dynamic Session Enhanced Key Protocol (DSEKP)—a lightweight, symmetric-only enhancement to the PSK framework that employs the HMAC-based Key Derivation Function (HKDF–SHA256) to generate ephemeral per-session encryption keys. DSEKP performs a single-round initialization handshake authenticated by an HMAC proof (InitProof) and derives fresh AES–GCM session keys from a combination of a device nonce, session counter, and timestamp. This design retains the simplicity of PSK while providing per-session key isolation (assuming the long-term secret remains uncompromised), replay protection, and dynamic key refresh without relying on public-key cryptography or certificate infrastructures.
@@
    Lightweight symmetric session protocol: DSEKP extends PSK with HKDF-based per-session key derivation and HMAC authentication, providing per-session key isolation and replay protection without public-key operations.
@@ 2 Related Work
The security of Internet of Things (IoT) communication remains an active research field focused on balancing computational efficiency with cryptographic robustness. While asymmetric standards such as TLS 1.3 and DTLS 1.3 provide confidentiality and mutual authentication through public-key exchanges, their multi-round handshakes and certificate management overheads management are challenging for constrained microcontrollers such as the ESP32, ATmega, or STM32 families [padmavathi2025security, kavitha2024secure]. Consequently, lightweight symmetric models based on the Pre–Shared Key (PSK) paradigm have been widely adopted for protocols including MQTT and CoAP [amanlou2021lightweight, kaganurmath2025dlks]. This section traces the evolution from static PSK schemes toward dynamic or session-based key management approaches, emphasizing the gaps that motivate the proposed DSEKP framework.
@@ 2.1 PSK-Based Encryption in IoT Systems
PSK architectures rely on a single long-term symmetric key shared between device and broker to perform AES–GCM encryption with minimal processing overhead. Studies such as  [kim2020page],  [hasan2024survey] and  [sovyn2019comparison] have shown efficient resource utilization on ESP32 microcontrollers using PSK–AES modes. However, static key reuse introduces significant vulnerabilities: once the PSK is compromised, all past and future traffic becomes decryptable, violating forward secrecy. Replay protection in such systems typically depends on timestamp or counter validation, which can desynchronize during Wi–Fi disruptions or low-power sleep cycles [yu2024multi, farha2020timestamp]. Moreover, at scale, managing thousands of PSKs manually increases operational burden and risks inconsistent key distribution, leaving gaps exploitable by adversaries [savant2025secure].
@@ 2.2 Dynamic Session-Key and Rekeying Protocols
To mitigate static-key weaknesses, researchers have explored session-based and hybrid key-management protocols. Protocols such as DTLS 1.3 and EDHOC (Ephemeral Diffie–Hellman Over COSE) [pajkos2025esp32, krawczyk2016optls] achieve forward secrecy through elliptic-curve key exchanges, but these computations are prohibitively expensive for constrained MCUs [fedrecheski2024performance, astorga2022revisiting, restuccia2020low]. In contrast, symmetric-only methods refresh session keys using derivation functions without public-key operations. The HMAC-based Key Derivation Function (HKDF) is particularly suited for IoT systems, as it expands a static secret using per-session entropy sources [shahidinejad2024all, akshatha2023mqtt, kaganurmath2025enabling]. For instance,  [rastoceanu2022blockchain, karmous2024hybrid] employed HKDF–SHA256 for MQTT rekeying but required a centralized key server to manage nonces.  [hakeem2021key, pinto2016hash, li2024hash] shown hash-chain–based key evolution to enhance secrecy, yet synchronization overhead limited scalability. Efforts such as DTLS [pittoli2016dtls, park2014lightweight] reduce handshake size through message compression but still depend on asymmetric initialization, leaving open the need for a purely symmetric, session-oriented protocol.
@@ 2.3 Comparative Gaps and Motivation
Comparative analyses consistently reveal a trade-off between cryptographic strength and computational feasibility.  [suarez2017practical] reported that ECDHE-based TLS increased energy consumption relative to PSK. Purely symmetric PSK models remain efficient but lack key agility and entropy mixing, making them susceptible to replay and traffic-analysis attacks [henriques2017using, yu2025forward]. Furthermore, existing HKDF-based schemes often depend on centralized infrastructures or omit practical hardware validation. Hence, a clear gap exists for a fully symmetric, self-contained, and experimentally validated session-key protocol that offers per-session key isolation (while the long-term secret remains secure), replay protection, and low overhead—without reliance on public-key infrastructures (PKI).
@@ 2.4 Contribution Context
The proposed Dynamic Session Enhanced Key Protocol (DSEKP) directly addresses these limitations. It preserves PSK’s lightweight efficiency while introducing HKDF-derived per-session keys and an HMAC-authenticated initialization handshake. Unlike prior HKDF or hash-chain approaches, DSEKP performs all derivation and verification steps locally on constrained devices without reliance on external key servers. Thus, it bridges the gap between the simplicity of static PSK and the key-agility of TLS, delivering a balanced, reproducible, and deployable solution for secure IoT–Edge communication.
@@ 3 Methodology and Framework
The proposed Dynamic Session Enhanced Key Protocol (DSEKP) extends the conventional Pre–Shared Key (PSK) model by introducing per-session dynamic key derivation and symmetric-only authentication using the HMAC-based Key Derivation Function (HKDF–SHA256). The protocol preserves the lightweight characteristics of PSK but replaces the static, reused key with ephemeral AES–GCM session keys derived from fresh entropy sources—device nonce, session counter, and timestamp—thereby providing per-session key isolation (assuming the long-term device secret remains uncompromised) and replay protection.. This section describes the system architecture, data communication workflow, protocol phases, algorithmic details, and integrated security mechanisms.
@@ 3.3 Protocol Workflow
    Session Termination / Rotation: Upon device reboot or timeout, a new session counter triggers a fresh HKDF derivation, providing per-session key isolation and automatic session cleanup on the edge node.
@@ 3.5 Security Operations
Per-Session Key Isolation HKDF entropy mixing Session keys remain independent while the long-term secret stays confidential (no forward secrecy if later compromised).
@@ 3.7 Summary
The DSEKP framework transforms the conventional static PSK architecture into a session-adaptive, stateless, and authenticated communication model for constrained IoT devices. By combining HKDF-based entropy mixing with HMAC-driven session authentication, DSEKP achieves lightweight key agility and per-session isolation without public-key operations and per-session key isolation with negligible additional cost in latency or bandwidth.
@@ 5.1 Descriptive Statistics
Interpretation and Justification: The results in Table 6 are derived directly from synchronized send–receive timestamps collected across 6,500 packet transmissions for each protocol, ensuring high statistical confidence. The 95 % confidence intervals (CI) confirm narrow uncertainty margins due to the large sample size and consistent network conditions. The observed 27 % increase in mean latency under DSEKP originates from the additional HKDF–SHA256 key derivation and HMAC verification performed once per session, introducing a small computational delay without affecting throughput stability. Similarly, the 10 % growth in mean payload corresponds to the inclusion of session metadata—namely the 2–byte session counter and 14–byte HMAC proof—required for per-session key isolation and authentication. Despite these expected overheads, both throughput and reliability remain nearly identical between PSK and DSEKP, confirming that the proposed dynamic rekeying mechanism introduces negligible runtime or bandwidth penalty. These results verify that DSEKP preserves the lightweight characteristics of PSK while enhancing security and timing determinism in constrained IoT environments.
@@ 5.4 Payload and Overhead Analysis
Figure 7: Payload size comparison between PSK and DSEKP. The $\approx$ 10 % increase corresponds to additional session–metadata fields required for dynamic authentication. This minor overhead enables per-session key isolation and replay protection without significant bandwidth penalty.
@@ 5.5 Security–Performance Trade–off Summary
Interpretation. DSEKP exhibits a moderate latency increase of approximately 27 % and a payload growth of about 10 % compared with PSK, while preserving identical throughput (1 pps) and near–perfect reliability (>>99.8 %). Although the latency difference is statistically significant (very low pp–values), the effect size remains moderate (Cohen’s $d\approx-0.49$), confirming that the overhead is small and within operational limits for lightweight IoT telemetry. Overall, DSEKP strengthens cryptographic resilience—providing per-session key isolation (assuming the long-term secret remains uncompromised) and replay protection—with negligible runtime penalty, offering a practical balance between performance and security for IoT–Edge deployments.
@@ 5.7 Interpretation of Results
The overall findings confirm that DSEKP effectively strengthens PSK–based security while preserving the lightweight characteristics required in IoT networks. The modest latency increase remains well within acceptable bounds for low–frequency telemetry ($\leq$ 1 Hz). DSEKP’s symmetric and stateless architecture allows seamless deployment on resource–constrained microcontrollers and edge gateways without modifying existing MQTT infrastructures. In summary, DSEKP delivers modern cryptographic agility—per-session key isolation, replay protection, and session isolation—at negligible computational and bandwidth cost, establishing it as a practical upgrade path for PSK systems in real IoT–Edge environments.
@@ 6.1 Security Improvements Over PSK
DSEKP derives a unique AES–GCM key for every session using the HMAC-based Key Derivation Function (HKDF–SHA256), mixing device entropy sources—a 12-byte random nonce (DevNonce), 2-byte session counter (SessCtr), and 4-byte timestamp (TT)—with a long-term secret and salt. This ensures that compromise of one session key cannot expose previous or future sessions, thereby providing per-session key isolation (assuming the long-term device secret remains uncompromised) without asymmetric cryptography.
@@ 6.2 Performance and Scalability Considerations
Despite introducing dynamic key derivation and per-session authentication, DSEKP’s runtime performance remains near-identical to static PSK. As shown in Figures 46, the mean latency increase of 27 % and 10 % payload growth are modest relative to the gains in per-session key isolation and replay protection, remaining within typical Wi–Fi variance for low-frequency telemetry (1–2 Hz).
@@ 6.3 Comparative Perspective
When positioned against standard IoT security frameworks such as DTLS 1.3, EDHOC, or LAKE, DSEKP provides a pragmatic middle ground between performance and cryptographic assurance. It delivers session-key agility with only symmetric operations using only symmetric primitives—achieving more than 80 % lower computational cost than DTLS while maintaining comparable latency to PSK.
@@ 6.6 Summary of Discussion
Overall, DSEKP transforms static PSK schemes into secure, adaptive, and stateless architectures suitable for modern IoT–Edge ecosystems. It provides measurable improvements in confidentiality, integrity, and provides per-session key isolation while maintaining sub-second latency and constant throughput. By combining HKDF-based entropy mixing with symmetric HMAC authentication, DSEKP represents a practical and scalable step toward trustworthy, autonomous, and energy-efficient IoT communication.
@@ 7 Conclusion and Future Work
This paper presented the Dynamic Session Enhanced Key Protocol (DSEKP), a lightweight and symmetric-only security framework that strengthens conventional Pre–Shared Key (PSK) encryption for Internet of Things (IoT) edge environments. By deriving fresh AES–GCM session keys through HKDF–SHA256 and authenticating session initialization via an HMAC-based proof, DSEKP achieves per-session key isolation (assuming the long-term device secret remains uncompromised), replay protection, and stateless key management—without relying on public-key cryptography.
@@
The proposed protocol effectively bridges the gap between PSK simplicity and lightweight key agility and per-session isolation without public-key operations, providing a deployable and scalable security solution for IoT–Edge communication. Its stateless, symmetric design supports seamless scaling to thousands of devices while maintaining strong per-session cryptographic isolation. In doing so, DSEKP establishes a practical foundation for trust-based and distributed edge security architectures in next-generation networks.
@@ Acknowledgments
Data Availability
The PSK baseline benchmark dataset [rg5z-ge26-25] and the proposed DSEKP dataset [3vw8-3a13-25] used in this study are publicly archived on IEEE DataPort. These resources include encrypted and decrypted MQTT Dataset logs to support reproducibility.